# Residual Diffusion Models for Variable-Rate Joint Source Channel Coding of MIMO CSI

### v1

@@ IV Residual Diffusion for CSI Enhancement
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{Z}_{0},\mathbf{T}}\left[\frac{\bar{\alpha}_{\mathbf{T}}}{1-\bar{\alpha}_{\mathbf{T}}}\left\|\mathbf{Z}_{0}-f_{\text{den}}(\mathbf{Z}_{\mathbf{T}},\hat{H}_{d},\mathbf{T};\theta_{\text{den}})\right\|^{2}\right],$ (11)
@@
IV-A Training
We propose a two-stage training strategy to optimize performance while maintaining computational efficiency. In the first stage, the autoencoder model is trained using the MSE loss (2) between the input CSI matrix and the estimated CSI matrix. Due to the relatively small number of parameters and faster convergence, the computational overhead for Stage 1 is minimal. The detailed training procedure for this stage is outlined in Alg. 1.
@@ IV-A Training
2:     Sample $H_{d}$ from the channel distribution
3:     $\mathbf{s}=f_{\text{enc}}(H_{d};\theta_{\text{enc}})$
4:     $\hat{H}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}})$
5:     $\mathcal{L}_{\text{MSE}}=\|H_{d}-\hat{H}_{d}\|^{2}$
6:     Adam$(\theta_{\text{enc}},\theta_{\text{dec}},\mathcal{L}_{\text{MSE}})$
@@
2:     Sample $H_{d}$ form the channel distribution
3:     $\mathbf{s}=f_{\text{enc}}(H_{d};\theta_{\text{enc}})$
4:     $\hat{H}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}});~{}~{}\mathbf{r}=\hat{H}_{d}-\mathbf{z}_{0}$
@@
6:     Adam$(\theta_{\text{den}},\mathcal{L}_{\text{diff}})$
@@
V Experimental Setup and Results
V-A Baselines and comparison.
@@ V Experimental Setup and Results
To evaluate our diffusion-based approach, we benchmark against ADJSCC [27], a recent state-of-the-art non-linear transform method. ADJSCC employs a neural network to transform CSI information from the spatial frequency domain to a low-dimensional representation—bypassing the traditional IFFT approach—before further compression via a secondary network. We also include the deep JSCC variant of CSINet+ [39], another widely recognized benchmark in CSI compression literature. To maintain fairness in comparison, we utilize identical datasets and training protocols across all evaluated models.
@@
We implement two variants of the ADJSCC baseline: (1) the original architecture as presented in [27], and (2) a parameter-scaled version that matches the complexity of our proposed RD-JSCC model, by increasing the layers and channel dimensions. Additionally, we examine a supervised variant of RD-JSCC, where both the autoencoder and U-Net components are trained end-to-end using an MSE loss function. Our primary contribution, RD-JSCC—which integrates an autoencoder with diffusion-based U-Net training—demonstrates performance improvements of an order of magnitude compared to [27] at equivalent NMSE targets.
V-B Dataset
@@
We evaluate our method on the COST2100 outdoor dataset [50], which is widely adopted due to its realistic and complex channel characteristics. The dataset provides $10^{5}$ training samples and $2\times 10^{4}$ test samples. It is important to note that the original spatial-frequency domain representations are not publicly available; instead, we utilize the provided $32\times 32$ cropped complex channel matrices in the angular-delay domain. Consequently, we omit the non-linear transformation module used in ADJSCC and retain only the inner encoder-decoder components that operate directly on the $32\times 32$ angular-delay domain complex inputs. Since the input dimensions of the inner module match those of the COST2100 cropped channels, this setup enables a fair and consistent comparison. All evaluations are reported in terms of NMSE measured in the angular-delay domain.
[…]
V-C Model configuration
@@
[…]
V-D Hyperparameters for training
@@
As described in Sec. IV, the model is trained in two stages. In the first stage, the autoencoder model is trained for $N_{\text{train}}=10^{5}$ iterations using the MSE loss (2). In the second stage, the diffusion-based U-Net is trained for $N_{\text{train}}=10^{6}$ iterations. The coarse estimate produced by the autoencoder serves as the initialization point for reverse diffusion, with denoising performed over $T=20$ steps during training. However, during inference, we can use a 2-step denoising for the reverse diffusion, with a small penalty in performance. This is primarily enabled by training the diffusion model using x-prediction instead of Gaussian noise, thus making the inference 10×10$\times$10 × faster. Performance is evaluated using the NMSE, which is given as $\mathbb{E}\left[{\|\mathbf{z}-\hat{\mathbf{z}}\|}/{\|\mathbf{z}\|}\right]$ for the ground truth $\mathbf{z}$ and reconstruction $\hat{\mathbf{z}}$.
V-E Results
@@
Our evaluation begins with the exact architectures from ADJSCC [27] and CSINet+ [39], which demonstrate notably poor performance. When scaling these architectures by incorporating additional intermediate layers and increasing convolutional channels, performance improves marginally, but all schemes still saturate at an NMSE greater than $-4$ dB. This reveals that conventional autoencoder approaches trained with supervised learning objectives exhibit poor scaling characteristics relative to model size and fail to adequately capture the underlying channel characteristics despite substantial parameter counts.
We then developed a hybrid architecture combining the original autoencoder model from [27] with a U-Net backbone for denoising. The U-Net backbone has been widely employed in image processing for enhanced denoising, deblurring, and, more recently, image generation. Our hybrid model, termed U-Net based JSCC, was trained with the same supervised training objective as [27], minimizing the end-to-end MSE. Remarkably, despite having a similar parameter count as the larger variants of ADJSCC and CSINet+, the U-Net-based JSCC shows significant performance improvements, achieving an NMSE lower than $-9$ dB. This highlights the importance of architectures that scale effectively with increasing model size.
Finally, we trained the U-Net-based JSCC model with a diffusion objective instead of the supervised objective. This approach, which we refer to as RD-JSCC, achieves the best performance among all evaluated schemes. During decoding at the receiver, the autoencoder head first produces a coarse estimate of the CSI. The U-Net model then initializes the reverse diffusion with this coarse estimate and refines the CSI at each denoising step, iteratively feeding the improved estimate back into the U-Net until completing the specified number of denoising steps. While our default configuration uses 20 denoising steps, we found that to minimize inference latency, a 2-step reverse diffusion performs remarkably well with only a negligible NMSE penalty, achieving an NMSE below $-12$ dB.
@@
Diffusion-based modeling demonstrates a clear advantage in capturing complex channel distributions, particularly when conventional supervised methods struggle to represent such distributions effectively.
Figure 4: RD-JSCC achieves an order-of-magnitude improvement in performance over state-of-the-art deep JSCC baselines on the COST2100 outdoor dataset for feedback bandwidth of $k=16$.
V-F NMSE Performance Translation to BLER
While NMSE is the widely adopted metric for reporting CSI compression performance, block error rate (BLER) provides a more direct measure of practical system reliability. We therefore evaluate the uncoded BLER performance to assess how NMSE improvements translate to error rate reductions. For the COST2100 dataset, we construct SF domain data from the available AD domain measurements using fast Fourier transform (FFT), assuming lossless SF to AD transformation in the original dataset. As shown in Fig. 5, both methods achieve similar BLER performance at low SNR values. However, in the high-SNR regime, the U-Net based approach saturates around $10^{-4}$ BLER, while the diffusion-based RD-JSCC continues improving and reaches BLER $\approx 10^{-6}$ at 5 dB uplink SNR. This performance difference correlates with the underlying reconstruction quality: U-Net based JSCC achieves NMSE $\approx-9.2$ dB while RD-JSCC achieves NMSE $\approx-12.2$ dB, at a SNR of 5 dB in the uplink.
@@ V-F NMSE Performance Translation to BLER
For the COST2100-outdoor dataset, at a uplink SNR of 5 dB, the 3 dB improvement in NMSE of the CSI translates to approximately two orders of magnitude improvement in BLER, demonstrating that the NMSE gains from diffusion-based approach translates to reduced error floor in high-SNR conditions for improved reliability.
@@
Figure 5: Uncoded BLER performance comparison between U-Net based JSCC and RD-JSCC approaches. The diffusion-based method shows continued improvement at high SNR, achieving BLER $\approx 10^{-6}$ at 5 dB while the U-Net approach saturates around $10^{-4}$.
V-G Computational Complexity

### v2

@@ IV Residual Diffusion for CSI Enhancement
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{Z}_{0},\mathbf{T}}\left[\frac{\bar{\alpha}_{\mathbf{T}}}{1-\bar{\alpha}_{\mathbf{T}}}\left\|\mathbf{Z}_{0}-f_{\text{den}}(\mathbf{Z}_{\mathbf{T}},\hat{\bm{H}}_{d},\mathbf{T};\theta_{\text{den}})\right\|^{2}\right],$ (11)
@@
Model Training. We propose a two-stage training strategy to optimize performance while maintaining computational efficiency. In the first stage, the autoencoder model is trained using the MSE loss (2) between the input CSI matrix and the estimated CSI matrix. Due to the relatively small number of parameters and faster convergence, the computational overhead for Stage 1 is minimal. The detailed training procedure for this stage is outlined in Alg. 1.
@@
2:  Sample $\bm{h}_{d}$ from the channel distribution
3:  $\mathbf{s}=f_{\text{enc}}(\bm{h}_{d};\theta_{\text{enc}})$
4:  $\hat{\bm{h}}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}})$
5:  $\mathcal{L}_{\text{MSE}}=\|\bm{h}_{d}-\hat{\bm{h}}_{d}\|^{2}$
[…]
@@
2:  Sample $\bm{h}_{d}$ form the channel distribution
3:  $\mathbf{s}=f_{\text{enc}}(\bm{h}_{d};\theta_{\text{enc}})$
4:  $\hat{\bm{h}}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}});~~\mathbf{r}=\hat{\bm{h}}_{d}-\mathbf{z}_{0}$
@@
[…]
@@
V Variable Rate Compression
In practical wireless communication systems, the available uplink bandwidth for CSI feedback varies dynamically due to factors such as performance requirements and network congestion. Traditionally, deep learning-based compression methods address multiple compression rates by training separate models for each rate. However, this approach imposes significant overhead on resource-constrained user equipment (UE). To mitigate this, recent research has focused on developing single-model solutions that can handle multiple compression rates.
For instance, CSI-PPPNet [10] introduces a one-sided architecture that supports arbitrary compression ratios using a single deep learning model. This design significantly reduces the number of parameters required at the UE, enhancing efficiency and simplifying deployment. Similarly, the Variable-Rate Code (VRC) approach dynamically allocates feedback bits based on estimated distortion, optimizing the trade-off between compression rate and reconstruction quality [24]. Additionally, the Variable Code Size Autoencoder (VCSA) framework adjusts codeword lengths during training, enabling flexible compression without retraining for different rates [47] but requires perfect transmission of the codeword indices, which is a disadvantage of vector quantization. Instead of a codebook-based approach, continuous-valued latents with scalar quantization naturally handle noisy transmission by mapping directly to subcarriers with power constraints.
Matryoshka Representation Learning. Introduced in [28], Matryoshka Representation Learning(MRL) is a representation learning framework designed to support multiple compression rates within a single model. The key idea behind MRL is to learn a single hierarchical representation that can be truncated to different lengths, yielding compressed representations of varying fidelity. This is achieved by partitioning the learned representation vector into nested subsets, such that each prefix subset corresponds to a progressively higher compression rate. Importantly, MRL enforces joint optimization of all nested representations during training, enabling the model to perform well across a broad spectrum of rate-distortion trade-offs without the need for retraining or maintaining multiple models. This design allows for efficient and adaptive deployment in scenarios with dynamic bandwidth or storage constraints, while maintaining representational consistency and strong generalization across different rates.
$\mathcal{L}_{\text{MRL}}=\sum_{i=1}^{m}\lambda_{i}\|\bm{H}_{d}-\hat{\bm{H}}_{d}^{(k_{i})}\|^{2},$ (12)
where $\hat{\bm{H}}_{d}^{(k_{i})}$ denotes the reconstructed CSI from the compressed representation of dimension $k_{i}$, and $\lambda_{i}$ are weighting factors that balance the importance across different rates. We choose $\lambda_{i}=\nicefrac{{1}}{{i}}$ by default, assigning lower weights to loss from higher-compression rates.
In Fig. 6, we present results for the model optimized across four feedback bandwidths $k\in\{8,16,24,32\}$, corresponding to compression rates $\rho\in\{\nicefrac{{1}}{{128}},\nicefrac{{1}}{{64}},\nicefrac{{1}}{{42}},\nicefrac{{1}}{{32}}\}$. The encoder architecture compresses CSI to a maximum complex-valued latent dimension of $32$, with the MRL training objective (Eq. 12) enabling nested prefix truncation to generate latents of dimensions $24$, $16$, and $8$. Crucially, while we explicitly optimize for four discrete compression levels, the nested structure inherently supports all intermediate rates, specifically, all integer bandwidths $k\in\{8,9,\ldots,32\}$, yielding $25$ distinct compression rates from a single unified model. During inference, the encoder adaptively selects the latent dimension based on available bandwidth, achieving an efficient rate-distortion trade-off without model switching overhead.
VI Experimental Setup and Results
VI-A Baselines and comparison.
@@ VI Experimental Setup and Results
To evaluate our diffusion-based approach, we benchmark against ADJSCC [54], a recent state-of-the-art non-linear transform method. ADJSCC employs a neural network to transform CSI data from the spatial-frequency domain into a low-dimensional representation, thereby bypassing the traditional IFFT approach, before further compression via a secondary network. We also include the deep JSCC variant of CSINet+ [14], another widely recognized benchmark in CSI compression literature. To ensure a fair comparison, we use identical datasets and training protocols across all evaluated models.
@@
We implement two variants of the ADJSCC baseline: (1) the original architecture as presented in [54], and (2) a parameter-scaled version that matches the complexity of our proposed RD-JSCC model, by increasing the layers and channel dimensions. Additionally, we examine a supervised variant of RD-JSCC in which both the autoencoder and U-Net components are trained end-to-end with an MSE loss. Our primary contribution, RD-JSCC, which integrates an autoencoder with diffusion-based U-Net training, demonstrates performance improvements of an order of magnitude compared to [54] at equivalent NMSE targets.
Architectural Differences. To ensure fair comparison across architectures, we maintain identical encoder-decoder designs for ADJSCC, CSINet+, and RD-JSCC, with key distinctions: (1) ADJSCC includes an analog transformation module that maps spatial-frequency domain CSI to a learned representation before compression, (2) ADJSCC and RD-JSCC incorporate SNR-adaptive feature scaling during encoding and decoding, while CSINet+ does not. Our RD-JSCC framework extends the base autoencoder architecture (ADJSCC without the transformation module) by adding a U-Net-based diffusion denoising module at the decoder. This modular design isolates the contribution of diffusion-based refinement from other architectural choices.
Although parameter matching via layer and dimension adjustments may not represent optimal baseline configurations, the order of magnitude performance difference when our approach outperforms the best non-diffusion baseline, as seen in Fig. 5, highlights fundamental limitations of existing supervised learning methods for joint source channel coding of CSI in complex channels. These substantial gains, even under capacity-constrained conditions, validate our architectural innovations beyond parameter-count considerations.
VI-B Dataset
@@
We evaluate our method on the COST2100 outdoor dataset [35], which is widely used for its realistic and complex channel characteristics. The dataset provides $10^{5}$ training samples and $2\times 10^{4}$ test samples. It is important to note that the original spatial-frequency domain representations are not publicly available; instead, we utilize the provided $32\times 32$ cropped complex channel matrices in the angular-delay domain. Consequently, we omit the non-linear transformation module used in ADJSCC and retain only the inner encoder-decoder components that operate directly on the $32\times 32$ angular-delay domain complex inputs. Since the input dimensions of the inner module match those of the COST2100 cropped channels, this setup enables a fair and consistent comparison. Following standard preprocessing practices, we apply min-max normalization to scale the complex channel matrices to $[0,1]$ by dividing by twice the maximum absolute value and adding 0.5. All evaluations are reported in terms of NMSE measured in the angular-delay domain.
Since there is no associated uplink channel for the COST2100 outdoor dataset, we generate uplink channel realizations using QuaDRiGa [21], adhering to the 3GPP TR 38.901 channel specification [49], with an uplink carrier frequency of $5.4~\text{GHz}$. We adopt the simulation configuration outlined in [54] and assume a dominant line-of-sight (LOS) component. The base station (BS) is positioned at the center of a $20~\text{m}\times 20~\text{m}$ area and is equipped with a uniform linear array (ULA) comprising $N_{\mathrm{t}}=32$ omnidirectional elements spaced at half-wavelength intervals. The user equipment (UE) employs a single omnidirectional antenna ($N_{\mathrm{r}}=1$). The antenna heights are set to $3\,\text{m}$ at the BS and $1.5\,\text{m}$ at the UE. The simulation includes $N_{c}=32$ subcarriers in the uplink transmission. A sampled channel realization in SF and AD domains for the 3GPP Indoor dataset is provided in Fig. 4, to explicitly show the sparse nature of the channel in the AD domain and the ability to truncate the channel without losing any significant information in the SF domain.
Figure 4: CSI preprocessing for 3GPP Indoor dataset. The spatial-frequency channel undergoes 2D IFFT transformation to the angular-delay domain, revealing sparsity that enables truncation to $32\times 32$ dominant coefficients. Reconstruction via 2D FFT achieves near-lossless recovery of the original channel.
VI-C Model configuration
@@
Diffusion. The architecture for the diffusion denoising network is based on [55] and [25]. We choose an initial embedding width of $64$ channels and dimension multipliers $\{1,\,2,\,3,\,4\}$ for the successive down-sampling and mirrored up-sampling stages. Skip connections link each down-sampling block to its symmetric up-sampling counterpart, preserving high-resolution features throughout the reverse diffusion trajectory. The U-Net outputs a refined estimate of the CSI matrix that is fed back, iteratively refining the coarse estimate, based on the number of steps used in reverse diffusion.
VI-D Hyperparameters for training
@@
As described in Sec. IV, the model is trained in two stages. In the first stage, the autoencoder model is trained for $N_{\text{train}}=10^{5}$ iterations using the MSE loss (2). In the second stage, the diffusion-based U-Net is trained for $N_{\text{train}}=10^{6}$ iterations. The coarse estimate produced by the autoencoder serves as the initialization point for reverse diffusion, with denoising performed over $T=20$ steps during training. However, during inference, we can use a 2-step denoising approach for reverse diffusion, with a small performance penalty. This is primarily enabled by training the diffusion model using x-prediction instead of Gaussian noise, thus making the inference $10\times$ faster. Performance is evaluated using the NMSE, which is given as $\mathbb{E}\left[{\|\mathbf{z}-\hat{\mathbf{z}}\|}/{\|\mathbf{z}\|}\right]$ for the ground truth $\mathbf{z}$ and reconstruction $\hat{\mathbf{z}}$.
Further, regarding the noise scheduling, our derivation in Equation (7) adopts deterministic DDIM sampling ($\sigma_{t}=0$) and a linear residual trajectory assumption. This design is motivated by two key factors for CSI feedback:
  • Reconstruction Consistency: Unlike generative tasks, CSI recovery requires high fidelity. Deterministic sampling eliminates stochastic variance, ensuring a unique, optimal reconstruction for each compressed latent.
  • Low-Latency Robustness: In few-step regime required for faster inference, deterministic sampling significantly outperforms stochastic methods (e.g., DDPM) which require many steps ($T\gg 50$) to average out noise [46].
VI-E Results
Figure 5: RD-JSCC achieves an order-of-magnitude improvement over state-of-the-art deep JSCC baselines on the COST2100 outdoor dataset for $k=16$. While traditional supervised learning approaches fail under complex channel distributions, our diffusion-based iterative refinement successfully models the underlying statistics.
@@ VI-E Results
Our evaluation begins with the exact architectures from ADJSCC [54] and CSINet+ [14], which demonstrate notably poor performance. When scaling these architectures by incorporating additional intermediate layers and increasing the number of convolutional channels, performance improves only marginally, yet all schemes still saturate at an NMSE greater than -4 dB. This reveals that conventional autoencoder approaches trained with supervised learning objectives exhibit poor scaling characteristics relative to model size and fail to adequately capture the underlying channel characteristics despite substantial parameter counts.
We then developed a hybrid architecture combining the original autoencoder model from [54] with a U-Net backbone for denoising. The U-Net backbone has been widely employed in image processing for enhanced denoising, deblurring, and, more recently, image generation. Our hybrid model, termed U-Net-based JSCC, was trained with the same supervised objective as [54], namely, minimizing end-to-end MSE. Remarkably, despite having a parameter count comparable to that of the larger variants of ADJSCC and CSINet+, the U-Net-based JSCC achieves significant performance improvements, achieving an NMSE below -9 dB. This highlights the importance of architectures that scale effectively with increasing model size.
Finally, we trained the U-Net-based JSCC model with a diffusion objective rather than a supervised objective. This approach, which we refer to as RD-JSCC, achieves the best performance among all evaluated schemes. During decoding at the receiver, the autoencoder head first produces a coarse estimate of the CSI. The U-Net model then initializes reverse diffusion with this coarse estimate and refines the CSI at each denoising step, iteratively feeding the improved estimate back into the U-Net until the specified number of denoising steps is completed. While our default configuration uses 20 denoising steps, we found that a 2-step reverse diffusion minimizes inference latency while incurring only a negligible NMSE penalty, achieving an NMSE below -12 dB. Notably, 1-step diffusion exhibits significant performance degradation.
@@
Diffusion-based modeling demonstrates a clear advantage in capturing statistically complex channel distributions that conventional supervised methods struggle to represent effectively. This advantage stems from the diffusion framework’s ability to learn high-dimensional probability distributions through iterative refinement.
Variable-Rate Compression. Further, we now consider the performance of RD-JSCC under the MRL objective to simultaneously optimize the performance for multiple rates, as shown in Fig. 6. The MRL model exhibits a rate-dependent performance characteristic: at lower bandwidths ($k\leq 16$), it achieves superior reconstruction compared to rate-specific models, while at higher bandwidths ($k\geq 24$), dedicated models show marginal gains. This behavior is expected, as the MRL objective optimizes average performance across the compression-rate spectrum (Eq. 12), trading slight suboptimality at individual rates for the practical advantages of supporting $25$ rates with a single model and eliminating the storage/switching overhead of maintaining multiple rate-specific architectures.
Figure 6: Optimizing for multiple feedback bandwidths: Using a single model with MRL objective, RD-JSCC dynamically adjusts the latent dimension based on available uplink bandwidth, achieving efficient rate-distortion trade-offs across multiple feedback rates.
VI-F NMSE Performance Translation to BLER
While NMSE is widely used to report CSI compression performance, block error rate (BLER) provides a more direct measure of practical system reliability. We therefore evaluate the uncoded BLER performance to assess how NMSE improvements translate to error rate reductions. For the COST2100 dataset, we construct SF-domain data from the available AD-domain measurements using the fast Fourier transform (FFT), assuming a lossless SF-to-AD transformation in the original dataset. As shown in Fig. 7, both methods achieve similar BLER performance at low SNR values. However, in the high-SNR regime, the U-Net-based approach saturates around $10^{-4}$ BLER, while the diffusion-based RD-JSCC continues improving and reaches BLER $\approx 10^{-6}$ at 5 dB uplink SNR. This performance difference correlates with the underlying reconstruction quality: U-Net-based JSCC achieves NMSE $\approx-9.2$ dB while RD-JSCC achieves NMSE $\approx-12.2$ dB, at an SNR of 5 dB in the uplink. It is interesting to note that the diffusion-based approach underperforms at low SNRs because it relies on the inductive bias of the learned channel distribution rather than the heavily corrupted initial estimate, generating statistically plausible but potentially inaccurate reconstructions, a known limitation of diffusion models under high noise conditions. Conversely, the supervised U-Net maintains stable performance even at low SNRs, while diffusion-based refinement excels at high SNRs where sufficiently accurate initial estimates enable effective iterative enhancement.
@@ VI-F NMSE Performance Translation to BLER
For the COST2100-outdoor dataset, at an uplink SNR of 5 dB, a 3 dB improvement in CSI NMSE corresponds to approximately two orders of magnitude improvement in BLER, demonstrating that NMSE gains from the diffusion-based approach translate into a reduced error floor at high SNR, thereby improving reliability.
@@
Figure 7: Uncoded BLER comparison for COST2100 shows that RD-JSCC continues improving at high SNR while U-Net-based JSCC saturates. The diffusion-based approach achieves BLER $\approx 10^{-6}$ at 5 dB, demonstrating two orders of magnitude improvement over the U-Net’s saturation at $10^{-4}$.
VI-G Computational Complexity